find the dimensions of a rectangle whose length is 3 more than twice its width and has a perimeter of 30 centimeters
Answer by user_dude2008(1862)

2W+2L=30
2W+2(2W+3)=30
2W+4W+6=30
6W+6=30
6W=30-6
6W=24
W=24/6
W=4
Width: 4
L=2W+3=2(4)+3=8+3=11
Length: 11